Hello and Happy Tuesday:)
After my first order of operations, parenthesis, would I consider 5(6)and 12(3)an exponent problem or a multiplication problem. I copied how I solved, but feel as if I am missing something.
Mathematical Problem of my own creation
300 - 5(3+3) = 306 - 12(3)
First I would solve the problem in parenthesis to simplify then rewrite the equation.
300 - 5(6) = 306 - 12(3)
Next I would solve all my multiplication problems to simplify then rewrite the equation.
300- 30 = 306 - 36
Finally I would solve all my subtraction problems to come up with the solution to the problem.
270 = 270

Found 2 solutions by ewatrrr, Theo:
Answer by ewatrrr(24785)

 
Hi,
Yes.
"PEMDAS", ( "Please Excuse My Dear Aunt Sally")
Order: "Parentheses, Exponents, Multiplication and Division, and Addition and Subtraction".

Answer by Theo(13342)

it's a multiplication problem.
looks like you solved it correctly.
exponent problem would be something like 5^2 or 5*5 which is the equivalent of 5^2
something like 5*3^2 would be an exponent problem.
you would do the exponentiation first and then do the multiplication.
5*3^2 = 5*9 = 45
if you did the multiplication first you would be wrong.
5*3^2 = 15^2 = 225 would be wrong.
however, if you surrounded the 5*3 with parentheses so it looked like this:
(5*3)^2, then multiplying 5*3 first would be appropriate.